jigglypuff314 (jigglypuff314):

Evaluate the limit: lim x->(pi/6)^+ (2x^2+5x-3)/(2-csc(x))

jigglypuff314 (jigglypuff314):

I had gotten \[\lim_{x \rightarrow \frac{ \pi }{ 6 }^{+}}\frac{ 2x ^{2}+5x-3 }{ 2-\csc x } \rightarrow \infty\]

hartnn (hartnn):

yes, correct. because the denominator is positive,

jigglypuff314 (jigglypuff314):

okay thank you :)

hartnn (hartnn):

and for x->pi/6^- , it would be -infinity welcome ^_^